OpenStudy (haleyelizabeth2017):

What is the equation, in standard form, of a parabola that contains the following points? (–2, –20), (0, –4), (4, –20) Please help me and then at end after I get answer can you show all the work we did?

OpenStudy (here_to_help15):

@hero @mathgeek27 @camerondoherty @amistre64

OpenStudy (here_to_help15):

@e.mccormick

OpenStudy (jdoe0001):

have you covered system of equations yet?

myininaya (myininaya):

We can help you discover the answer for yourself. :) So first of all do you know what an equation for a parabola looks like?

OpenStudy (haleyelizabeth2017):

yes. \[y=ax^2+bx+c\]

OpenStudy (haleyelizabeth2017):

right?

myininaya (myininaya):

right

myininaya (myininaya):

you are given 3 points...

myininaya (myininaya):

This should give us a system of equations to solve.

myininaya (myininaya):

\[f(x)=ax^2+bx+c \\ \\ f(-2)=-20 \\ f(0)=-4 \\ f(4)=-20\]

OpenStudy (haleyelizabeth2017):

yes. I have those 3 down....one second :)

OpenStudy (haleyelizabeth2017):

\[−20=(−2)^2a+(−2)b+c.....(−4)=(0)^2a+(0)b+c.......-20=(4)2a+(4)b+c\] the ......separates the equations...

OpenStudy (haleyelizabeth2017):

\[−20=(−2)^2a+(−2)b+c ..(−4)=(0)^2a+(0)b+c..−20=(4)2a+(4)b+c\]

myininaya (myininaya):

\[4a-2b+c=-20 \\ c=-4 \\ 16a+4b+c=-20\] Just to make sure I'm seeing things right You have this correct?

OpenStudy (haleyelizabeth2017):

I don't know :(

OpenStudy (haleyelizabeth2017):

yes :) I just hadn't simplified it yet....

myininaya (myininaya):

ok so you have c=-4 so I suggest going to the other two equations and replacing c with -4

OpenStudy (haleyelizabeth2017):

okay......one second

myininaya (myininaya):

You now have a system of two equations with 2 variables

OpenStudy (haleyelizabeth2017):

so: 4a-2b-4=-20 16a+4b-4=-20

OpenStudy (haleyelizabeth2017):

and then do I add 4 to both sides?

OpenStudy (haleyelizabeth2017):

so: 4a-2b=-16 16a+4b=-16 8a-4b=-32 16a+4b=-16 24a=-48 a=-2

OpenStudy (haleyelizabeth2017):

4(-2)-2b-4=-20 -12-2b=-20 -2b=-8 b=4 a=-2 b=4 c=-4 correct?

myininaya (myininaya):

one sec while i check this

myininaya (myininaya):

\[4a-2b-4=-20 \\ 16a+4b-4=-20 \\ \\ 4a-2b=-16 \\16a+4b=-16 \\ 2a-b=-8 \\ 4a+b=-4 \\ 6a=-12 \\ a=-2 \\ 2(-2)-b=-8 \\-4-b=-8 \\ -b=-4 \\b=4\] looks good

OpenStudy (haleyelizabeth2017):

thank you so much!!!

myininaya (myininaya):

You did all the work so congrats @haleyelizabeth2017

myininaya (myininaya):

You could pass a test on this now.

OpenStudy (haleyelizabeth2017):

:)

myininaya (myininaya):

so you can go back to our y=ax^2+bx+c and replace a with -2, b with 4, and c with -4 and you will have your quadratic in standard form

OpenStudy (haleyelizabeth2017):

okay :) so \(y=-2x^2+4b-4\) correct?

myininaya (myininaya):

y=-2x^2+4x-4

myininaya (myininaya):

I think you meant x instead of b :p

OpenStudy (haleyelizabeth2017):

oops! I meant x yes :)

OpenStudy (here_to_help15):

hi

OpenStudy (here_to_help15):

your in good hands @haleyelizabeth2017 she is a master mind

OpenStudy (haleyelizabeth2017):

heeeyyyyy I haveta finish my math really quick :)

OpenStudy (here_to_help15):

@myininaya is

OpenStudy (here_to_help15):

okie

myininaya (myininaya):

I wouldn't say I'm a master mind. Also haley is really smart too so she herself has capable hands.

OpenStudy (here_to_help15):

You both have a capable minds

myininaya (myininaya):

I think you do too here_to_help15.

OpenStudy (haleyelizabeth2017):

heehee thanks! I wish we could turn it in with the dropbox so I can do it on paper! ugh but no, we have to use stupid workpad :'( It sucks :(

OpenStudy (here_to_help15):

ikr

OpenStudy (haleyelizabeth2017):

I very much dislike workpad if you can't tell :/

myininaya (myininaya):

I don't think I ever heard of workpad.

OpenStudy (haleyelizabeth2017):

You're not missing out on anything! :P

myininaya (myininaya):

Is that a math symbols type of thing?

myininaya (myininaya):

Like I mean you can type everything out in math symbols

OpenStudy (haleyelizabeth2017):

its where we have to turn in our math assignments.......

myininaya (myininaya):

hmm.. well good luck with the workpad thing. It sucks it isn't user friendly. When we are all old I bet it will be so simple to turn homework in online.

OpenStudy (haleyelizabeth2017):

okay :) thanks!

OpenStudy (haleyelizabeth2017):

it was simple until they took down the dropbox :(

OpenStudy (here_to_help15):

hey